OpenStudy (laddiusmaximus):

find an equation of the tangent line to y=2sin(x) x=3pi/4

OpenStudy (laddiusmaximus):

I know that 2sinx become 2cosx

OpenStudy (laddiusmaximus):

so is that 2cos(3pi/4)? which becomes \[-\sqrt{4}\]?

OpenStudy (anonymous):

\[2\cos(\frac{3\pi}{4})=2\times -\frac{\sqrt{2}}{2}=-\sqrt{2}\]

OpenStudy (laddiusmaximus):

so does that become y-\[\sqrt{4}\]=\[-\sqrt{4}\](x-3\[\Pi\]/4

OpenStudy (laddiusmaximus):

y-sqrt4=sqrt4(x-3pi/4)

OpenStudy (laddiusmaximus):

right?

OpenStudy (accessdenied):

satellite was correct with the value being negative sqrt(2)

OpenStudy (laddiusmaximus):

wait what?

OpenStudy (laddiusmaximus):

oh got it.

OpenStudy (laddiusmaximus):